OpenStudy (anonymous):

A container in the form of a triangular prism contains water at a depth of 6 inches when pointed downwards. It can contain 12 inches water. The height of prism is 18 inches and its base is 8 inches. Find the volume of water and the wet part of container. Please help.... :)

OpenStudy (anonymous):

Surace area of base = 8(4√3)/2 = 16 √3 Volume of water = 16 √3 (18) = 288 √3 cubic inch Wet part = 16 √3 + 18(8 + 8 + 8) = 16 √3 + 432 square inch

OpenStudy (anonymous):

@liliegirl

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

Nope...

OpenStudy (phi):

 Here is a picture of what it looks like. I assume the "base" is the top part of the triangle. they say the prism can contain 12 inches of water, which I assume means its "height" is 12 inches. It only has 6 inches of water, so the water goes half-way up.